Journalist Report – December 24th

Crew 201 Journalist Report – 24DEC2018

The Martian Chronicles

[SOL 8] – It’s beginning to look a lot like Christmas

"It’s the most wonderful time of the year.
There’ll be much mistletoing,
and hearts will be glowing,
when loved ones are near.
It’s the most wonderful time of the year"

MDRS Crew 201 – MEx-1

Back on Earth, we celebrate Christmas Eve with our families and loved ones. Traditionally we get together and have dinner, share presents and remember the best times of this year. We remember those who passed before us, and how their memories are still with us.

These moments we cherish the most, feel warmth in our hearts and try to make them last forever. In Mars, it could not be any different, we got together to make a tree out of cardboard, paint a few decorations (Thank you Commander Ilaria!) and bring memories all together.

Work today, aside from our holiday, there were several pre-dinner adventures. Some level of success has been shown in the Green Hab experiment where local soil is being used as a testbed for future crop growing. So far, 3 of the 10 different soil combinations have germinated radish. Can you imagine? We have our very first Mars Organically grown Radish! Well, maybe not us but Crew 202 will which to us is quite a fantastic moment.

An extended EVA was performed to the southeast of the Hab where Juan Carlos and Chilpo were working on artificial vision and camera testing for the rover. Traverse to the location was performed on the rover Curiosity but the road was complex and they missed their exit by a few kilometers and had to turn back to the North Pinto Hills.

After completing a walk of around 2 miles some strange footprints were found along the road. Strange noises were heard around the EVA location and the footprints, strange for an inhabited planet, seemed to be from a known earth species that travel with an old man in a red suit. Santa Claus and Rudolph, was that you?

Jokes aside, our Christmas Eve has been fantastic. We miss our families, but in our new home, the warmth of the season is felt by every crewmember.

From the Crew-201 stationed in Mars, we want to wish you all ¡Feliz Navidad!

MEx-1 continuará informando.

Genaro Grajeda, MDRS Crew 201 Co-Journalist

EVA Report – December 24th

Crew 201 EVA Report 24-Dec-2018

EVA #7

Author: Carlos Mariscal, XO

Purpose of EVA: To do artificial vision and camera testing in various terrains in order to determine algorithm performance in environments with changes of ground color and rock sizes and to look for appropriate locations for rover prototype testing.

Start time: 11:00 h

End time: 15:00 h

Narrative: We started our journey with a really good mood, drove to the southeast using the rover Curiosity and did a couple of stops in Robert’s Rock Garden and Zubrin’s Head, ideal landscapes for some pictures to train our artificial vision algorithms, then we drove to find the Greenstone road, but it’s quite hard to find, we got lost and almost got to the highway, but we turned back and looked carefully, left the rover and started walking, heading to North Pinto Hills.

Amazing landscapes and pics! We walked around 3 km and climbed one of the hills; on our way back to the rover we found several animal footprints, maybe a feline, elks and mice, quite scary at first, but something nice to see. We got safe to the rover and started our way back and Something really random happened; we found a truck in the way, a family (dad and two children) were excited to see us and asked a picture of the kids with us. Quite funny.

Finally, we got home, safe and sound, with some cool stories for the rest of the crew and lots of photos, both to remember and keep on working.

Destination: North Pinto Hills, with stops at Robert’s Rock Garden and Zubrin’s Head.

Coordinates (use UTM NAD27 CONUS): N 4248782.15, E 520960.16

César Serrano, Engineer

Carlos Mariscal, XO

Road(s) and routes per MDRS Map: Driving to south-east on Cow Dung Road 0110 then walking to north-east Greenstone Road and finally east to North Pinto Hills.

Mode of travel: Rover Curiosity

Sol Summary – December 24th

Sol: 9

Summary Title: Christmas from Mars

Author’s name: Tania Robles, Commander

Mission Status: Mission objectives are going fine; every crew member is working on each project. However today we dedicated more time to cook and clean the Hab to celebrate Christmas.

Sol Activity Summary: Last night we went to bed late because we saw The Martian movie. We woke up at 8 am and had our daily routine with breakfast and briefings. It seems that we are not losing much weight, just around 2 kilograms per person until today, Sol 9.

Two members went out to an EVA were they found life in the form on animals and a couple of humans who asked for a picture to our astronauts. After they came back, we worked in our Christmas tree (nice artistic skills, Crew!). By now we are finishing our special dinner with lasagna.

Happy Christmas up to Earth!

Look Ahead Plan: Tomorrow we will work individually in our projects but also record some videos to Mexico so we can send them on our way back to Earth. Also, we will have a 3D Printing class and have dinner with the MDRS Director.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: All good.

EVA: XO and Crew Engineer went to North Pinto Hills.

Reports to be filed: Operations Report
EVA Report
Greenhab Report
Astronomy Report
Journalist Report
Pictures of the day

Support Requested: A lot of discussions went into what Star Wars movies were directed or produced by George Lucas. Can you help us with the list of the 8 episodes, please?

Operations Report – December 24th


Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None

Generator (hours run): 17:06 yesterday until around 8: 15. Today started at 17:08

Solar— SOC 77%
Diesel Reading – Around 71%
Propane Reading – 74%
Ethanol Free Gasoline – N/A

Water (Auxillary tank) – 0 gallons
Water (static tank) – High Level, around 215 -gallons
Axillary to Static tank transfer– no
Gallons transferred: 0
Water in GreenHab – Medium level (47%)
Water (loft) – Static to Loft Pump used – yes
Water Meter: 01394271

Toilet tank emptied: Yes

Deimos rover used: no, still not functional
Hours: 0
Beginning charge: –
Ending charge: –
Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging X :

Spirit rover used: NO
Hours: 63
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Opportunity rover used: NO
Hours: 44.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Curiosity rover used: YES
Hours: 61.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Notes on rovers: We used curiosity for today’s EVA.
# Hours the ATVs were Used today: Nothing to report.
Notes on ATVs: Nothing to report.

HabCar used and why, where? Not used.
CrewCar used and why, where? Not used.

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit 4 still has not enough air flow.

Summary of GreenHab operations: – Nothing to report.

Summary of ScienceDome operations: – Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Musk Solar observatory – Robotic Observatory – Nominal.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: MERRY CHRISTMAS!

Thank you!!!

Greenhab Report – December 24th

Crew 201 Greenhab Report 24-DEC-2018
Greenhab Officer: Walter Calles

Environmental Control:
Cooling w/ ambient air (1.5hrs)

40% Shade Cloth on

Average Temperatures:
Low: 19°C
High: 38°C

Hours of Supplemental Light: 5

Daily Water Usage of Crops: 12 gallons

Water in Blue Tank (lbs): 48%

Times of Watering for Crops: 1230

Changes to crops:
Re-planted salad beds (mix and blend) started to grow
Chives looking good, first sprouts growing
Peas growing good
Baby greens showing excellent progress
Tomatoes growing slow
Excellent progress on lettuce
Fennel looking better
Carrots growing slow but fine

Some harvesting for Christmas Eve dinner today. Plants are in good shape. Spinaches and onions still with no changes, hopefully, they will grow properly.

Experiment Results:
Date: SOL4 (SOL9 overall)
First results observed!
Control Pot: No results yet
Soil Sample Type One
@30%: First radish leaves
@20: First radish leaves
Soil Sample Type 2
@10%: First radish leaves

Salad Blend: 153g
Radish: 87g
Cilantro: 25g
Basil: 7g
Oregano: 1g

Support/Supplies Needed: none

Merry Christmas!

Copyright © The Mars Society. All rights reserved. | Main Site